OpenStudy (anonymous):

A football is thrown at an angle of 50◦ above the horizontal. Assume the ball is throw and received at the same height. To throw a 59.6 m pass, what must be the initial speed of the ball? The acceleration due to gravity is 9.8 m/s2 . Answer in units of m/s

OpenStudy (agent0smith):

Hopefully you've seen the range equation \[\Large R = \frac{ v_o ^2 \sin 2 \theta }{ g }\] R is distance, theta is angle, v0 is initial velocity note this equation ONLY APPLIES FOR NO CHANGE IN HEIGHT

OpenStudy (anonymous):

sin2(50) does that equal to sin(100) I've always wondered that

OpenStudy (agent0smith):

Yes. It's sin(2 theta) not sin2 * theta, since that wouldn't make much sense anyway

OpenStudy (anonymous):

Oh well I got 24.3524 m/s

OpenStudy (agent0smith):

Well then, good job :)

OpenStudy (anonymous):

Thank you :)